"a wallet contains $5 bills and $10 bills. There are 15 bills in the wallet with a total value of $120. find the number of $5 bills and $10 bills" Im having trouble figuring this problem out can u show me how to set it up and solve?
Answer by stanbon(75887)

"a wallet contains $5 bills and $10 bills. There are 15 bills in the wallet with a total value of $120. find the number of $5 bills and $10 bills"
-------------
Quantity Equation: f + t = 15 bills
Value Equation::::5f +10t = 120 dollars
--------------------------------------------
Multiply thru the Quantity Eqution by 5 to get:
5f + 5t = 75
---------------
Subtract that Eq. from the Value equation and solve for "t":
5t = 45
t = 9 ($ of ten dollar bills)
-----
Since f+t = 15, f +9 = 15, and f = 6 (# of 5 dollar bills)
================================================================
Cheers,
Stan H.